Solve the linear system in the next question using the Substitution Method. Describe the steps.
2x + 3y = -4
x – 2y = 5

Answer by checkley77(12844)

2x + 3y = -4
x – 2y = 5 or x=2y+5
now substitute (2y+5) for x in the first equation & solve for y.
2(2y+5)+3y=-4
4y+10+3y=-4
7y=-4-10
7y=-14
y=-14/7
y=-2 ans.
Now replace y by -2 in the second equation & solve for x.
x-2(-2)=5
x+4=5
x=5-4
x=1 ans.
Proof:
2*1+3*-2=-4
2-6=-4
-4=-4